# Transcoders Find Interpretable LLM Feature Circuits

### v1

@@ 4.2 Blind case study: reverse-engineering a feature
To understand the utility of transcoders for circuit analysis, we carried out nine blind case studies, where we randomly selected individual transcoder features in a ninth-layer (of 12) GPT2-small transcoder and used circuit analysis to form a hypothesis about the semantics of the feature—without looking at the text of examples that cause the feature to activate. In blind case studies, we use a combination of input-invariant and input-dependent information to allow us to evaluate transcoders as a tool to infer model behavior with minimal prompt information. We believe that if our circuit analysis tools can cleanly disentangle input-dependent and input-invariant information, then we can better understand the extent to which their insights are likely to generalise and allow us to predict out-of-distribution behavior, a key goal for mechanistic interpretability. The “rules of the game” for blind case studies are as follows:
@@
Current-token features. From token 37, we found strong contributions from tc0[16632]@37 and tc0[9188]@37. Input-invariant de-embeddings of these layer 0 features revealed that they primarily activate on variants of ;, causing us to hypothesize that token 37 contributed to the feature by virtue of being a semicolon. Another feature which contributed strongly through the current token, tc6[11831], showed a similar pattern. Among the top input-invariant connections from layer 0 transcoder features to tc6[11831], we once again found the same semicolon features tc0[16632] and tc0[9188].
@@ 4.3.1 Initial investigation
Next, we investigated which MLP10 transcoder features were most important on a variety of prompts, and how their activations are mediated by attention head 1 in layer 9. Following the original work, we generated all 100 prompts of the form “The war lasted from 17YY to 17”, where YY denotes a two-digit number. We found that the MLP10 features with the highest variance in activations over this set of prompts also had top input-dependent connections from MLP0 features through attention head 1 in layer 9 whose top de-embeddings were two-digit numbers. We then turned to look at the MLP0 features with the top input-invariant connections to these MLP10 features, as mediated by attention head 1 in layer 9. Looking at the top de-embedding tokens in turn for these MLP0 features gives us input-invariant information about which tokens in the input are most important for causing the MLP10 features to activate.
As it turned out, almost always among the top five de-embedding tokens for these MLP10 features were two-digit numbers—indicating that two-digit number tokens were among the most important for causing these MLP10 features to fire out of all tokens in the model’s vocabulary, without reference to any specific input. This positive result was somewhat unexpected, given that there are only 100 two-digit number tokens in the model’s vocabulary of over 50k tokens.
We then used direct logit attribution (DLA) [9] to look at the effect of each transcoder feature on the predicted logits of each YY token in the model’s vocabulary. These results, along with normalized de-embedding scores (details in App. G) for each YY token in the model’s vocabulary, can be seen in Figure 4. We see that the de-embedding scores are highest for YY tokens where years following them are boosted and years preceding them are inhibited.
Figure 4: For the three MLP10 transcoder features with the highest activation variance over the “greater-than” dataset, and for every possible YY token, we plot the DLA (the extent to which the feature boosts the output probability of YY) and the de-embedding score (an input-invariant measurement of how much YY causes the feature to fire).
@@ 4.3.2 Comparison with neuronal approach
Next, we compared the transcoder approach to the neuronal approach to see whether transcoders give a sparser description of the circuit than MLP neurons do. To do this, we computed the highest-variance layer 10 transcoder features and MLP10 neurons. Then, for $1\leq k\leq 65$, we zero-ablated all but the top $k$ features in the transcoder/neurons in MLP10 and measured how this affected the model’s performance according to the mean probability difference metric presented in the original paper. We also evaluated the original model with respect to this metric, along with the model when MLP10 is replaced with the full transcoder.
@@
The results are shown in the left half of Figure 5. For fewer than 24 features, the transcoder approach outperforms the neuronal approach; its performance drops sharply, however, around this point. Further investigation revealed that tc10[5315], the 24th-highest-variance transcoder feature, was responsible for this drop in performance. The DLA for this feature is plotted in the right half of Figure 5. Notice how, in contrast with the three highest-variance transcoder features, tc10[5315] displays a flatter DLA, boosting all tokens equally. This might explain why it contributes to poor performance. To account for this, note that Figure 5 also demonstrates the performance of the transcoder when this “bad feature” is removed.
@@ 5 Related work
Circuit analysis is a common framework for exploring model internals [7, 9, 10]. A number of approaches exist to find circuits and meaningful components in models, including causal approaches [32], automated circuit discovery [33], and sparse probing [16]. Causal methods include activation patching [36, 37, 38], attribution patching [39, 40], and path patching [41, 8]. Much circuit analysis work has focused on attention head circuits [42], including copying heads [9], induction heads [11], copy suppression [43], and successor heads [44]. Methods connecting circuit analysis to SAEs include He et al. [45], Batson et al. [46] and Marks et al. [20].
Sparse autoencoders have been used to disentangle model activations into interpretable features [17, 18, 19]. The development of SAEs was motivated by the theory of superposition in neural representations [47]. Since then, much recent work has focused on exploring and interpreting SAEs, and connecting them to preexisting mechanistic interpretability techniques. Notable contributions include tools for exploring SAE features, such as SAE lens [48]; applications of SAEs to attention sublayers [27]; scaling up SAEs to Claude 3 Sonnet [49], and improved SAE architectures [50].
Transcoders have been originally proposed as a variant of SAEs under the names “predicting future activations”  [23] and “MLP stretchers” [24]. More recently, concurrent work by Ge et al. [51] also investigated the use of transcoders (presented in their paper under the name of “skip-SAEs”) for performing circuit analysis, and applied them to understanding “bracketed text” features and the GPT2-small “indirect object identification” circuit first studied by Wang et al. [8]. Our contributions differ from those of  Ge et al. [51] in a number of ways, including our factorization of attributions into input-dependent and input-invariant terms, our transcoder interpretability analysis, our investigation of the greater-than circuit, and our blind case studies that attempt to reverse-engineer unseen features. Furthermore, we evaluate the sparsity and accuracy of transcoders across multiple hyperparameters and larger models, whereas Ge et al. [51] evaluate multiple layers in GPT2-small at a single hyperparameter. Despite these differences, we find it exciting to see that the general utility of transcoders for mechanistic interpretability is being more broadly recognized.
@@ Appendix B Compute details
The most compute-intensive parts of the research presented in this work were training the SAEs and transcoders used in Section 3.2.2, along with the set of GPT2-small transcoders used in Sections 4.2 and 4.3. Training all of these SAEs and transcoders involved GPUs. The SAEs and transcoders from Section 3.2.2 were trained on an internal cluster using an A100 GPU with 80 GB of VRAM. The VRAM used by each training run ranged from approximately 16 GB for the GPT2-small runs to approximately 60 GB for the Pythia-1.4B runs. The time taken by each training run ranged from approximately 30 minutes for the GPT2-small transcoders/SAEs to approximately 3.5 hours for the Pythia-1.4B runs.
@@
In addition to these training runs, our case studies were carried out on internal cluster nodes with GPUs. These case studies used no more than 6 GB of VRAM. The total amount of compute used during each case study is variable (depending on how in-depth one wants to investigate a case study), but is de minimis in comparison to the training runs. The same goes for the computation of top activating examples used in Section 3.2.1.
@@ D.4 Recursing on a single computational path
To do this, we will extend the intuition presented in Equation 21 and previously discussed by Dunefsky and Cohan [53], which is to propagate our feature vector backwards through the computational path. [^1] The similarity to backpropagation is not coincidental, as it can be shown that the method about to be described computes the “input-times-gradient” attribution often used in the explanability literature. Starting at the end of the computational path, for each node in the computational path, we compute the attribution of the node towards causing the current feature vector to activate; we then compute a new feature vector, and repeat the process using the preceding node and this new feature vector.
@@
There is one caveat, however, that must be noted. Before every sublayer in the transformer architectures considered in this paper (that is, before every MLP sublayer and attention sublayer), there is a LayerNorm nonlinearity. Neel Nanda [39] provides intuition that LayerNorm nonlinearities can be approximated as a linear transformation that scales its input by a constant; Dunefsky and Cohan [53] provide further theoretical motivation and empirical results suggesting that this is reasonable. We follow this approach in our circuit analysis by multiplying each $\mathbf{f}$ feature vector by the appropriate LayerNorm “scaling constant” (which is empirically estimated by taking the ratio of the norm of the pre-LayerNorm activation vector to the post-LayerNorm activation vector).
@@ D.5 Full circuit-finding algorithm
Output
@@
Initialize $\mathcal{P}\leftarrow\{[(\mathbf{f^{\prime}},l^{\prime},t^{\prime},a^{\prime})]\}$triangleright $\mathcal{P}$ will be our working set of computational paths. Each computational path is a list of feature vectors paired with their attributions.
Initialize $\mathcal{P}_{out}\leftarrow\{\}$triangleright This will contain our output
@@
     Initialize $\mathcal{P}_{next}\leftarrow\{\}$triangleright This will contain the next iteration of computational paths
@@
         Initialize $\mathcal{A}\leftarrow\{\}$triangleright The set of attributions of all lower-layer features
@@
Output
     $\mathcal{G}=(\mathcal{V},\mathcal{E})$ A computational graph formed from the paths of $\mathcal{P}$.
Initialize $\mathcal{S}\leftarrow\{\}$triangleright A set of already-seen computational path prefixes, to prevent us from double-counting attributions
Initialize $\mathcal{V}\leftarrow\{\}$triangleright A dictionary mapping nodes to their attributions
Initialize ${}\mathcal{E}\leftarrow\{\}$triangleright A dictionary mapping edges (node pairs) to their attributions
@@
              Set $n_{parent}\leftarrow P[i-1]$, $n_{child}\leftarrow P[i]$triangleright Earlier-layer nodes come later in the computational paths returned by Algorithm 1
@@
Appendix E Details on Section 3.2.2 SAE/transcoder training
In this section, we provide details on the hyperparameters used to train the SAEs and transcoders evaluated in Section 3.2.2.
@@
Appendix F Details on Section 3.2.1
The transcoder used in the interpretability comparison was the Pythia-410M layer 15 transcoder trained with $\lambda_{1}$ sparsity coefficient $5.5\times 10^{-5}$ from Section 3.2.2. The SAE used in the comparison was a Pythia-410M layer 15 SAE trained on MLP inputs with $\lambda_{1}=7.0\times 10^{-5}$. We used an SAE trained on MLP inputs rather than one trained on MLP outputs (as in § 3.2.2) because the interpretability comparison involves looking at which examples cause features to activate. This, in turn, is wholly determined by the encoder feature vectors. Because the transcoder’s encoder feature vectors live in the MLP input space, it is thus most valid to compare the transcoder to an SAE whose encoder feature vectors also live in the MLP input space.
@@
To obtain the de-embedding scores shown in Figures 4 and 5, the following method was used. First, we used the method presented in Appendix D.3 to determine which MLP0 transcoder features had the highest input-invariant connections to the given MLP10 transcoder feature through attention head 1 in layer 9. Specifically, for MLP0 transcoder feature $i$ and MLP10 transcoder feature $j$, this attribution is given by $\left(\mathbf{f_{dec}^{(0,i)}}\right)^{T}\left(\mathbf{W_{OV}^{(9,1)}}\right)^{T}\mathbf{f_{enc}^{(10,j)}}$. For each MLP10 transcoder feature, the top ten MLP0 transcoder features were considered. Then, for each MLP0 transcoder feature, the de-embedding score of each YY token for that MLP0 feature was computed. The total de-embedding score of each YY token for an MLP10 feature was computed as the sum of the de-embedding scores of that token over the top ten MLP0 features, with each de-embedding score weighted by the input-invariant attribution of the MLP0 feature. In Figures 4 and 5, the de-embedding scores were scaled and recentered in order to fit on the graph.
@@ H.1.1 Citation feature: tc8[355]
Input-independent investigation of tc6[21046]@35 revealed high contributions from tc0[16382] and tc0[5468]. tc0[16382] corresponded to tokens such as oglu, owski, and zyk; tc0[5468] corresponded to tokens such as Burnett, Hawkins, and MacDonald. Observing that all of these are (components of) surnames, we decided that token 35 was likely (part of) a surname.
@@ H.2.4 Four more restricted blind case studies
We present the results of four more restricted blind case studies in Table 3. In the interest of conserving space, only the results of these case studies are presented; the original Jupyter Notebooks in which the studies were carried out are available in our code, which can be found at https://github.com/jacobdunefsky/transcoder_circuits.

### v2

@@ 5.1 Blind case study: reverse-engineering a feature
To understand the utility of transcoders for circuit analysis, we carried out nine blind case studies, where we randomly selected individual transcoder features in a ninth-layer (of 12) GPT2-small transcoder and used circuit analysis to form a hypothesis about the semantics of the feature—without looking at the text of examples that cause the feature to activate. In blind case studies, we use a combination of input-invariant and input-dependent information to allow us to evaluate transcoders as a tool to infer model behavior with minimal prompt information. This better reflects a key goal of mechanistic interpretability: to be able to understand model behavior on unknown, unforeseen tasks.
In contrast, reverse-engineering a feature where one already has an idea of its behavior can introduce confirmation bias. For instance, looking at activation patterns prior to circuit analysis can predispose a researcher to seek out only circuits that corroborate their interpretation of these activation patterns, potentially ignoring circuits that reveal other information about the feature. Conversely, if the circuit analysis method is faulty and yields some explanations that are not reflected in the feature activations, then the researcher might ignore those spurious explanations and thus obtain an overly-positive assessment of the circuit analysis method. The “rules of the game” for blind case studies are that:
@@
Current-token features. From token 37, we found strong contributions from tc0[16632]@37 and tc0[9188]@37. Input-invariant de-embeddings of these layer 0 features revealed that they primarily activate on variants of ;, suggesting that token 37 contributed to the feature by virtue of being a semicolon. Another feature which contributed strongly through the current token, tc6[11831], showed a similar pattern. Among the top input-invariant connections from layer 0 transcoder features to tc6[11831], we once again found the same semicolon features tc0[16632] and tc0[9188].
@@ 5.2.1 Initial investigation
Next, we investigated which MLP10 transcoder features were most important on a variety of prompts, and how their activations are mediated by attention head 1 in layer 9. Following the original work, we generated all 100 prompts of the form “The war lasted from 17YY to 17”, where YY denotes a two-digit number. We found that the MLP10 features with the highest variance in activations over this set of prompts also had top input-dependent connections from MLP0 features through attention head 1 in layer 9 whose top de-embeddings were two-digit numbers. The top input-invariant connections from MLP0 features through attention head 1 in layer 9 to MLP10 features also had two-digit numbers among their top de-embedding tokens. This positive result was somewhat unexpected, given that there are only 100 two-digit number tokens in the model’s vocabulary of over 50k tokens.
We then used direct logit attribution (DLA) [15] to look at the effect of each transcoder feature on the predicted logits of each YY token in the model’s vocabulary. These results, along with de-embedding scores for each YY token, can be seen in Figure 5. The de-embeddings scores are highest for YY tokens where years following them are boosted and years preceding them are inhibited.
@@ 5.2.2 Comparison with neuronal approach
Next, we compared the transcoder approach to the neuronal approach to see whether transcoders give a sparser description of the circuit than MLP neurons do. To do this, we computed the 100 highest-variance layer 10 transcoder features and MLP10 neurons. Then, for $1\leq k\leq 100$, we zero-ablated all but the top $k$ features in the transcoder/neurons in MLP10 and measured how this affected the model’s performance according to the mean probability difference metric presented in the original paper. We also evaluated the original model with respect to this metric, along with the model when MLP10 is replaced with the full transcoder.
@@
The results are shown in the left half of Figure 4. For fewer than 24 features, the transcoder approach outperforms the neuronal approach; its performance drops sharply, however, around this point. Further investigation revealed that tc10[5315], the 24th-highest-variance transcoder feature, was responsible for this drop in performance. The DLA for this feature is plotted in the right half of Figure 4. Notice how, in contrast with the three highest-variance transcoder features, tc10[5315] displays a flatter DLA, boosting all tokens equally. This might explain why it contributes to poor performance. To account for this, note that the left half of Figure 4 also demonstrates the performance of the transcoder when this “bad feature” is removed.
@@
Figure 5: For the three MLP10 transcoder features with the highest activation variance over the “greater-than” dataset, and for every possible YY token, we plot the direct logit attribution (the extent to which the feature boosts the output probability of YY) and the de-embedding score (an input-invariant measurement of how much YY causes the feature to fire).
@@ 6 Related work
Circuit analysis is a common framework for exploring model internals [41, 15, 32]. A number of approaches exist to find circuits and meaningful components in models, including causal approaches [20], automated circuit discovery [11], and sparse probing [24]. Causal methods include activation patching [52, 49, 28], attribution patching [39, 30], and path patching [22, 50]. Much circuit analysis work has focused on attention head circuits [18], including copying heads [15], induction heads [42], copy suppression [35], and successor heads [23]. Methods connecting circuit analysis to SAEs include He et al. [27], Batson et al. [1] and Marks et al. [34]. Our recursive greedy circuit-finding approach was largely based on that of Dunefsky & Cohan [13].
Sparse autoencoders have been used to disentangle model activations into interpretable features [7, 12, 51]. The development of SAEs was motivated by the theory of superposition in neural representations [17]. Since then, much recent work has focused on exploring and interpreting SAEs, and connecting them to preexisting mechanistic interpretability techniques. Notable contributions include tools for exploring SAE features, such as SAE lens [5]; applications of SAEs to attention sublayers [29]; scaling up SAEs to Claude 3 Sonnet [48] and improved SAE architectures [45]. Transcoders have been previously proposed as a variant of SAEs under the names “predicting future activations”  [47] and “MLP stretchers” [31], but not explored in detail.
@@ Appendix B Compute details
The most compute-intensive parts of the research presented in this work were training the SAEs and transcoders used in Section 4.2, along with the set of GPT2-small transcoders used in Sections 5.1 and 5.2. Training all of these SAEs and transcoders involved GPUs. The SAEs and transcoders from Section 4.2 were trained on an internal cluster using an A100 GPU with 80 GB of VRAM. The VRAM used by each training run ranged from approximately 16 GB for the GPT2-small runs to approximately 60 GB for the Pythia-1.4B runs. The time taken by each training run ranged from approximately 30 minutes for the GPT2-small transcoders/SAEs to approximately 3.5 hours for the Pythia-1.4B runs.
@@
In addition to these training runs, our case studies were carried out on internal cluster nodes with GPUs. These case studies used no more than 6 GB of VRAM. The total amount of compute used during each case study is variable (depending on how in-depth one wants to investigate a case study), but is de minimis in comparison to the training runs. The same goes for the computation of top activating examples used in Section 4.1.
@@ D.4 Recursing on a single computational path
To do this, we will extend the intuition presented in Equation 23 and previously discussed by Dunefsky & Cohan [13], which is to propagate our feature vector backwards through the computational path. [^3] The similarity to backpropagation is not coincidental, as it can be shown that the method about to be described computes the “input-times-gradient” attribution often used in the explanability literature. Starting at the end of the computational path, for each node in the computational path, we compute the attribution of the node towards causing the current feature vector to activate; we then compute a new feature vector, and repeat the process using the preceding node and this new feature vector.
@@
There is one caveat, however, that must be noted. Before every sublayer in the transformer architectures considered in this paper (that is, before every MLP sublayer and attention sublayer), there is a LayerNorm nonlinearity. Neel Nanda [39] provides intuition that LayerNorm nonlinearities can be approximated as a linear transformation that scales its input by a constant; Dunefsky & Cohan [13] provide further theoretical motivation and empirical results suggesting that this is reasonable. We follow this approach in our circuit analysis by multiplying each $\mathbf{f}$ feature vector by the appropriate LayerNorm “scaling constant” (which is empirically estimated by taking the ratio of the norm of the pre-LayerNorm activation vector to the post-LayerNorm activation vector).
@@ D.5 Full circuit-finding algorithm
  The input prompt on which we will perform circuit analysis
  Output:
@@
  Initialize $\mathcal{P}\leftarrow\{[(\mathbf{f^{\prime}},l^{\prime},t^{\prime},a^{\prime})]\}$$\mathcal{P}$ will be our working set of computational paths. Each computational path is a list of feature vectors paired with their attributions. }
  Initialize $\mathcal{P}_{out}\leftarrow\{\}$This will contain our output}
  Run the model on the input prompt, caching all of its activations.
@@
     Initialize $\mathcal{P}_{next}\leftarrow\{\}$This will contain the next iteration of computational paths}
@@
        Initialize $\mathcal{A}\leftarrow\{\}$The set of attributions of all lower-layer features}
@@
  Output: $\mathcal{G}=(\mathcal{V},\mathcal{E})$ A computational graph formed from the paths of $\mathcal{P}$.
  Initialize $\mathcal{S}\leftarrow\{\}$A set of already-seen computational path prefixes, to prevent us from double-counting attributions}
  Initialize $\mathcal{V}\leftarrow\{\}$A dictionary mapping nodes to their attributions}
  Initialize ${}\mathcal{E}\leftarrow\{\}$A dictionary mapping edges (node pairs) to their attributions}
@@
           Set $n_{parent}\leftarrow P[i-1]$, $n_{child}\leftarrow P[i]$Earlier-layer nodes come later in the computational paths returned by Algorithm 1}
@@
Appendix E Details on Section 4.2 SAE/transcoder training
In this section, we provide details on the hyperparameters used to train the SAEs and transcoders evaluated in Section 4.2.
@@
Appendix F Details on Section 4.1
The transcoder used in the interpretability comparison was the Pythia-410M layer 15 transcoder trained with $\lambda_{1}$ sparsity coefficient $5.5\times 10^{-5}$ from Section 4.2. The SAE used in the comparison was a Pythia-410M layer 15 SAE trained on MLP inputs with $\lambda_{1}=7.0\times 10^{-5}$. We used an SAE trained on MLP inputs rather than one trained on MLP outputs (as in § 4.2) because the interpretability comparison involves looking at which examples cause features to activate. This, in turn, is wholly determined by the encoder feature vectors. Because the transcoder’s encoder feature vectors live in the MLP input space, it is thus most valid to compare the transcoder to an SAE whose encoder feature vectors also live in the MLP input space.
@@
To obtain the de-embedding scores shown in Figures 5 and 4, the following method was used. First, we used the method presented in Appendix D.3 to determine which MLP0 transcoder features had the highest input-invariant connections to the given MLP10 transcoder feature through attention head 1 in layer 9. Specifically, for MLP0 transcoder feature $i$ and MLP10 transcoder feature $j$, this attribution is given by $\left(\mathbf{f_{dec}^{(0,i)}}\right)^{T}\left(\mathbf{W_{OV}^{(9,1)}}\right)^{T}\mathbf{f_{enc}^{(10,j)}}$. For each MLP10 transcoder feature, the top ten MLP0 transcoder features were considered. Then, for each MLP0 transcoder feature, the de-embedding score of each YY token for that MLP0 feature was computed. The total de-embedding score of each YY token for an MLP10 feature was computed as the sum of the de-embedding scores of that token over the top ten MLP0 features, with each de-embedding score weighted by the input-invariant attribution of the MLP0 feature. In Figures 5 and 4, the de-embedding scores were scaled and recentered in order to fit on the graph.
@@ H.1.1 Citation feature: tc8[355]
Input-independent investigation of tc6[21046]@35 revealed high contributions from tc0[16382] and tc0[5468].  feat016382 corresponded to tokens such as oglu, owski, and zyk; tc0[5468] corresponded to tokens such as Burnett, Hawkins, and MacDonald. Observing that all of these are (components of) surnames, we decided that token 35 was likely (part of) a surname.
@@ H.2.4 Four more restricted blind case studies
We present the results of four more restricted blind case studies in Table 3. In the interest of conserving space, only the results of these case studies are presented. However, in the supplemental material attached to this submission, the original Jupyter Notebooks in which the case studies were carried out are provided.
@@
Table 4: The results of all blind case studies.
Feature Type Final hypothesis Actual interpretation Outcome
tc8[355] Blind Fires on semicolons in the context of academic citations Fires on semicolons in the context of academic citations Success
tc8[1447] Restricted Blind Single-token “All” feature Single-token “All” feature Success
tc8[6569] Restricted Blind Fires on prompts conveying “told/said in an interview” Fires on prompts conveying “told/said in an interview” Success
tc8[4911] Restricted Blind Fires on though or although in the beginning of a clause Fires on though or although in the beginning of a clause Success
tc8[6414] Restricted Blind Largely uninterpretable feature that sometimes fires on Cyrillic text Largely uninterpretable feature that sometimes fires on Cyrillic text Success
tc8[235] Blind Fires on past-tense synonyms of “to catch” Fires on forms of “caught” Mostly a success
tc8[2725] Restricted Blind Fires on phrases about not offering things or not providing things. (As a stretch: particularly in legalese context?) Fires on phrases about not offering things or not providing things, in general Mostly a success
tc8[479] Restricted Blind Fires on prompts resembling “with a total estimated…” A local context feature Failure
tc8[9030] Restricted Blind Fires on biology when in the context of being a subject of study Fires on scientific subjects of study like chemistry, psychology, biology, economics Failure